### v1

@@ 5.1 Gamma Distribution
$\pi({x})\propto\prod_{i=1}^{d}x_{i}^{\alpha_{i}-1}e^{-\beta_{i}x_{i}}=e^{-U({x})},$
where
$U({x})=\sum_{i=1}^{d}\beta_{i}x_{i}-(\alpha_{i}-1)\log(x_{i}).$
@@
This distribution is particularly challenging to simulate using Langevin dynamics since the gradient of $U$ is non-Lipschitz. One could follow the approach of [39] by introducing a regularization of $U$ leading to a gradient Lipschitz approximation as well as a reflection to deal with the positivity constraint. However, the introduction of regularization introduces bias and we thus expect this approach not to be competitive. We will hence only compare MZZS with various mirror Langevin algorithms for the barrier function $\psi$ defined by
$\psi(x)=\frac{1}{2}\|x\|^{2}-\sum_{i=1}^{d}\log x_{i}.$ (31)
The potential $V$ in the dual space defined in (11) is $L_{V}$-gradient Lipschitz with $L_{V}=\frac{1}{4}+\max_{1\leq i\leq d}\left[\frac{\alpha_{i}}{10}+\frac{\beta_{i}}{4}\right]$.
For simulation we set $d=1,\alpha=3$, and $\beta=10$. We use $5\times 10^{6}$ gradient evaluations to generate $5\times 10^{6}$ samples from MLAa, MLAm and MZZS. Note that in one dimension MZZS and MBPS coincide. For each set of samples we calculate the Wasserstein-1 error between the samples and true samples from the Gamma distribution. This is repeated $100$ times and Figure 4 gives box plots of these Wasserstein errors. We have included a box plot showing the error between copies of the true samples which we use to measure the Monte Carlo error coming from the estimation of the Wasserstein error.
Figure 4: Box plots of 100100 realisations of the $W^{1}$ error of MLAa, MLAm and MZZS using $5\times 10^{6}$ gradient evaluations.
@@ 5.2 Truncated Gaussian
Our objective is to sample from the truncated Gaussian distribution in dimension $d=10$ and defined on the open set ${K}=(0,5)\times(0,0.5)^{9}$ by $\pi(x)\propto e^{-U(x)}$ with $U(x)=\displaystyle\frac{1}{2}x^{\top}\Sigma^{-1}x$. The covariance $\Sigma$ of the target is defined by $\Sigma_{ij}=(1+|i-j|)^{-1}.$ We use the mirror map $\psi$ defined as
$\psi(x)=-\sum_{i=1}^{d}\log(x_{i}-a_{i})+\log(b_{i}-x_{i}),$ (32)
with $a_{i}=0,\ \text{for all}\ i,\ b_{1}=5,\ b_{2}=b_{3}=\ldots=b_{10}=0.5$. We compare both the performance of MZZS and MBPS to MLAa and MLAm. Furthermore, for the special case of the truncated Gaussian distribution, Wall Hamiltonian Monte Carlo (WHMC), ([40]) provides an exact way to sample and thus we consider the results obtained by it as the ground truth.
We plot in Figure 5 the box plots for the expectations of the first three components using the algorithms discussed above. In each case, we have used $10^{2}$ realisations of the algorithms. In the case of MLAa and MLAm, one step of the algorithm costs one gradient evaluation, while for MZZS and MBPS, generating a candidate inter-jump time $\tau^{*}$ costs one gradient evaluation. For MLAa, MLAm, MZZS and MBPS, each realisation uses $10^{7}$ gradient evaluations. Furthermore, for MZZS and MBPS once a skeleton has been constructed, we generate $10^{7}$ samples. The number of gradient evaluations used for WHMC is $10^{6}$ as this takes the same runtime as the other algorithms. For each realisation of the samplers we implement a $10\%$ burn in phase.
Figure 5: Boxplots of the expectations $\mathbb{E}[x_{1}],\ \mathbb{E}[x_{2}]$ and $\mathbb{E}[x_{3}]$ for the truncated Gaussian distribution in 10 dimensions, estimated using MLAa, WHMC, MLAm, MZZS and MBPS algorithms. The grey line in each box represent the true mean of the associated component.
As we can see in Figure 5, for the first component both MLAa and MLAm are biased, with MLAa having greater variance. There is no visible bias for the mirror PDMPs algorithms, as they agree with WHMC. In the case of the second and third components all the samplers behave in a similar unbiased manner.
5.3 Latent Dirichlet Allocation
A distribution frequently occuring in machine learning applications [41, 42], particularly in the area of text modeling [43] is the Dirichlet distribution. A typical problem in this area is the latent Dirichlet allocation (LDA). The aim of LDA is to find topics a document belongs to, based on the words in it.
@@ 5.3 Latent Dirichlet Allocation
$\pi({x}|{n})\propto\prod_{i=1}^{d}x_{i}^{n_{i}+\alpha_{i}-1},\quad{x}\in\Delta_{d}=\left\{x=(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}:\ \sum_{i=1}^{d}x_{i}=1,\ x_{i}>0\right\}.$
@@
The challenge in sampling from the Dirichlet distribution arises from the fact that the negative log density $U$ is not convex and the distribution is constrained to the simplex $\Delta_{d}$. The most interesting regime of the Dirichlet posterior occurs when it is sparse. Specifically, this means that most of the $n_{i}$’s are zero and only few of them are relatively large of the order of $\mathcal{O}(d)$ ([24]).
Note that the simplex $\Delta_{d}$ is a (d1)(d-1)-dimensional surface; hence in order to ensure integrability of the push-forward distribution, it is more convenient to work with $\mathbb{R}^{d-1}$ dimensional vectors. With this in mind, we use the following barrier function
$\psi(x)=\sum_{i=1}^{d-1}x_{i}\log x_{i}+\left(1-\sum_{i=1}^{d-1}x_{i}\right)\log\left(1-\sum_{i=1}^{d-1}x_{i}\right),\ \ x\in\Delta_{d}$ (33)
where we have used the equation of the simplex to reduce dimensionality by one.
For simulation, we create a synthetic dataset by generating $10^{4}$ data points from a $5$-dimensional multinomial distribution with a random probability vector $p$ (which is given in Section B.3). The parameters of the Dirichlet prior are $\alpha_{i}=0.1$ for all $i$. The dataset is then divided into $50$ batches of equal size.
@@
We evaluate the performance of MZZS, MBPS and the MZZ-SS algorithms compared to MLAa (12), MLAm (14) as well as their stochastic versions SMLAa and SMLAm. For MZZ-SS, as discussed in Section 3.4, we use a control variate with reference point being the mode of the push-forward $e^{-V}$. We will measure the cost of each algorithm in terms of epochs where an epoch denotes the equivalent cost of a full pass through the data. For MZZS, MBPS, MLAa and MLAm, one iteration involves one gradient evaluation, which corresponds to one epoch. On the other hand, MZZ-SS, SMLAa and SMLAm use a subset of the data at every iteration and therefore we can perform several iteration for the cost of one epoch. All the algorithms are run for $2.5\times 10^{5}$ epochs (More details on implementation in Appendix). For MZZS, MBPS and MZZ-SS, once the skeleton is constructed, we generate $2.5\times 10^{5}$ samples keeping track of the cost.
@@
Once the samples from each algorithms are obtained, we compute the running relative error in the standard deviation of the first component as a function of the number of epochs and the results are summarised in Figure 6. For the distribution $\pi$, the true standard deviation of the first component is $\sigma_{\text{true}}=4.26\times 10^{-3}.$ As we can observe in Figure 6, both MLAa and MLAm converge faster than MZZS and MBPS to their respective limiting behaviour but demonstrate significant bias. As expected, MZZS and MBPS have vanishing bias. SMLAa converges faster than MLAa but show higher levels of bias. In contrast, MZZ-SS converges faster than MZZS but remains unbiased. Since the constant $L_{V}^{L}$ does not exist (see details in Section B.3) we plot the error for different step sizes $\Delta t\in\{(L_{V})^{-1},(2L_{V})^{-1},(5L_{V})^{-1}\}$.
@@
Figure 6: A plot of the running relative error in the standard deviation of first component of the samples generated by MZZS, MBPS, MLAa, MLAm on the left and MZZS-$S$, SMLAa, SMLAm on the right
@@ 6 Conclusion and future work
In this paper we proposed a novel class of samplers, mirror PDMPs, which combine PDMPs with the concept of the mirror map. We demonstrated that such samplers benefit from the advantages of PDMPs, namely unbiasedness and the ability to perform exact subsampling while also respecting constraints on the probability distribution. These methods were compared to state of the art stochastic differential equation-based methods such as PLMC, MYULA, MLAa and MLAm. In each of our experiments we saw that MZZS and MBPS outperformed the SDE-based algorithms.
@@
In this paper we have concentrated on the most popular PDMPs, ZZS and BPS, however due to the nature of the mirror map it is often the case that the push-forward distribution on the dual space is heavy tailed and therefore it may be beneficial to consider other choices for the deterministic dynamics such as those introduced in [44].
Another direction for future work is to include the use of data driven mirror maps such as in [45], this could be combined with an adaptive approach to preconditioning [46]. A final future direction addresses a general with working with PDMPs which is to find sharp bounds for the switching rates, this has been addressed in [47, 48, 49, 50] and these techniques can be explored in the context of mirror PDMPs.
@@ Appendix B Calculating Lipschitz constants
In this section, we provide the derivation of the constants $L_{V},\ L_{V}^{S},\ L_{V}^{L}$ and $M_{\psi}$ which are used for the step size of the different problems encountered in Section 5.
@@
Computation of $L_{V}$. For the barrier function defined in (31), the Fenchel conjugate $\nabla\psi^{*}$ is defined by
@@
Computation of $M_{\psi}$. The barrier $\psi$ defined in (31) is 1-self concordant as the sum of self concordant functions [51, Theorem 5.1.1], hence $M_{\psi}=1$.
Computation of $L_{V}^{L}$. For all $x>0$ we have

### v2

@@ 5.1 Gamma Distribution
$\pi({x})\propto\prod_{i=1}^{d}x_{i}^{\alpha_{i}-1}e^{-\beta_{i}x_{i}}=e^{-U({x})},\quad U({x})=\sum_{i=1}^{d}\beta_{i}x_{i}-(\alpha_{i}-1)\log(x_{i}).$
@@
This distribution is particularly challenging to simulate using Langevin dynamics since the gradient of $U$ is non-Lipschitz. One could follow the approach of [Savvas2023] by introducing a regularization of $U$ leading to a gradient Lipschitz approximation as well as a reflection to deal with the positivity constraint. However, the introduction of regularization introduces bias and we thus expect this approach not to be competitive. We will hence only compare MZZS with various mirror Langevin for the barrier function $\psi$ defined in Table 1 for the positive orthant $(a_{i}=0,\ \forall i)$.
For the simulations we consider dimensions $d=1$ and $d=100$ with shape parameters $\alpha_{i}=3$, and rate parameters $\beta_{i}=10$ for all $i$. We use $5\times 10^{6}$ gradient evaluations to generate $5\times 10^{6}$ samples from MLAa, MLAm, MAMLA and MZZS. For each set of samples we calculate the Wasserstein-1 error between the empirical distribution of the first component and the corresponding true samples of the Gamma distribution. This is repeated $100$ times and Figure 4 gives box plots of these Wasserstein errors. We have also included a box plot showing the error between copies of the true samples which we use to measure the Monte Carlo error coming from the estimation of the Wasserstein error.
(a) $d=1$
(b) $d=100$
Figure 4: Box plots of $100$ realisations of the $W^{1}$ error of MLAa, MLAm, MAMLA and MZZS using $5\times 10^{6}$ gradient evaluations.
@@ 5.2 Truncated Gaussian
Our objective is to sample from the truncated Gaussian distribution in dimension $d=10$ and defined on the open set ${\mathcal{M}}=(0,5)\times(0,0.5)^{9}$ by $\pi(x)\propto e^{-U(x)}$ with $U(x)=\displaystyle\frac{1}{2}x^{\top}\Sigma^{-1}x$. The covariance $\Sigma$ of the target is defined by $\Sigma_{ij}=(1+|i-j|)^{-1}.$ We use the mirror map $\psi$ defined in Table 1 for the rectangle. We compare both the performance of MZZS and MBPS to MLAa, MLAm and MAMLA. Furthermore, for the special case of the truncated Gaussian distribution, Wall Hamiltonian Monte Carlo (WHMC), ([pakman2014exact]) provides an exact way to sample and thus we consider the results obtained by it as the ground truth.
We plot in Figure 5 the box plots for the expectations of the first three components using the algorithms discussed above. In each case, we have used $10^{2}$ realisations of the algorithms. In the case of MLAa, MLAm and MAMLA, one step of the algorithm costs one gradient evaluation, while for MZZS and MBPS, generating a candidate inter-jump time $\tau^{*}$ costs one gradient evaluation. For MLAa, MLAm, MAMLA, MZZS and MBPS, each realisation uses $10^{7}$ gradient evaluations. Furthermore, for MZZS and MBPS once a skeleton has been constructed, we generate $10^{7}$ samples. The number of gradient evaluations used for WHMC is $10^{6}$ as this takes the same runtime as the other algorithms. For each realisation of the samplers we implement a $10\%$ burn in phase.
Figure 5: Boxplots of the expectations $\mathbb{E}[x_{1}],\ \mathbb{E}[x_{2}]$ and $\mathbb{E}[x_{3}]$ for the truncated Gaussian distribution in 10 dimensions, estimated using MLAa, WHMC, MLAm, MAMLA, MZZS and MBPS algorithms. The grey line in each box represent the true mean of the associated component.
As we can see in Figure 5, for the first component both MLAa and MLAm are biased, with MLAa having greater variance. There is no visible bias for the mirror PDMPs and MAMLA, as they agree with WHMC. In the case of the second and third components all the samplers behave in a similar unbiased manner.
5.3 Bayesian Inference on Dirichlet posterior
A distribution frequently occurring in machine learning applications [mcauliffe2006nonparametric, fei2005bayesian], particularly in the area of text modeling [blei2003latent] is the Dirichlet distribution. A typical problem in this area is the latent Dirichlet allocation (LDA). The aim of LDA is to find topics a document belongs to, based on the words in it. In this section, we focus on sampling from a Dirichlet posterior which is a fundamental component of the full LDA model with the additional advantage that the true solution is known.
@@ 5.3 Bayesian Inference on Dirichlet posterior
$\pi({x}|{n})\propto\prod_{i=1}^{d}x_{i}^{n_{i}+\alpha_{i}-1},\quad{x}\in\Delta_{d}=\left\{x=(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}:\ \sum_{i=1}^{d}x_{i}=1,\ x_{i}\geq 0\right\}.$
@@
The challenge in sampling from the Dirichlet distribution arises from the fact that the negative log density $U$ is not convex and the distribution is constrained to the simplex $\Delta_{d}$. The most interesting regime of the Dirichlet posterior occurs when it is sparse. Specifically, this means that most of the $n_{i}$’s are zero and only few of them are relatively large of the order of $\mathcal{O}(d)$ ([hsieh2018mirrored]). Note that the simplex $\Delta_{d}$ is a (d1)(d-1)-dimensional surface; hence in order to ensure integrability of the push-forward distribution, it is more convenient to work with $\mathbb{R}^{d-1}$ dimensional vectors. We use the barrier function defined in Table 1 for the simplex.
@@
For simulation, we create a synthetic dataset by generating $10^{4}$ data points from a $5$-dimensional multinomial distribution with a random probability vector $p$ (which is given in Section B.5). The parameters of the Dirichlet prior are $\alpha_{i}=0.1$ for all $i$. We evaluate the performance of MZZS, MBPS and the MZZ-SS algorithms compared to MLAa (12), MLAm (14) and MAMLA as well as their stochastic versions SMLAa and SMLAm. For MZZ-SS, as discussed in Section 3.4, we use a control variate with reference point being the mode of the push-forward $e^{-V}$, while in addition we use mini-batches of size $200$. We will measure the cost of each algorithm in terms of epochs where an epoch denotes the equivalent cost of a full pass through the data. For MZZS, MBPS, MLAa, MLAm and MAMLA, one iteration involves one gradient evaluation, which corresponds to one epoch. On the other hand, MZZ-SS, SMLAa and SMLAm use a subset of the data at every iteration and therefore we can perform several iteration for the cost of one epoch. All the algorithms are run for $2.5\times 10^{5}$ epochs (More details on implementation in Appendix B.5). For MZZS, MBPS and MZZ-SS, once the skeleton is constructed, we generate $2.5\times 10^{5}$ samples keeping track of the cost.
@@
Once the samples from each algorithms are obtained, we compute the running relative error in the standard deviation of the first component as a function of the number of epochs and the results are summarised in Figure 6. For the distribution $\pi$, the true standard deviation of the first component is $\sigma_{\text{true}}=4.26\times 10^{-3}.$ As we can observe in Figure 6, both MLAa and MLAm converge faster than MZZS, MBPS and MAMLA to their respective limiting behaviour but demonstrate significant bias. As expected, MZZS and MBPS have vanishing bias and converge faster than MAMLA. SMLAa converges faster than MLAa but show higher levels of bias. In contrast, MZZ-SS converges faster than MZZS but remains unbiased. Since the constant $L_{V}^{L}$ does not exist (see details in Section B.5) we plot the error for both MLAm and SMLAm with step size $\Delta t=(10L_{V})^{-1}$.
@@
Figure 6: Plot of the running relative error in the standard deviation of the first component for samples generated by MZZS, MBPS, MLAa, MLAm,MAMLA (left) and by MZZS-$S$, SMLAa, SMLAm (right)
5.4 Bayesian reconstruction of quantized Gaussian processes
We consider a scenario in which we observe a function $f:\mathbb{R}\to\mathbb{R}$ through a quantization process. Let $f_{i}=f(x_{i}),\ i=1,\dots,n$ denote the true signal evaluated at given locations $x_{i}$. Instead of observing $f_{i}$ directly, we only observe a quantized version $y_{i}$ taking values in a finite set $\{q_{1},\dots,q_{N_{q}}\}$. Our goal is to reconstruct the values $f(x_{i})$ from these quantized observations. The quantization rule is given by
$y_{i}=q_{k}\quad\text{if}\quad z_{k}\leq f_{i}<z_{k+1}$ (31)
where $\{z_{k}\}_{k=1}^{N_{z}}$ are the quantization thresholds. Denoting by $y=(y_{1},\dots,y_{n})$ and $f=(f_{1},\dots,f_{n})$, we aim to sample from the posterior distribution $\pi(f|y)$. We impose a Gaussian prior $\pi_{0}(f)$ on $f$ with covariance $\Sigma$, i.e. $f\sim\mathcal{N}(0,\Sigma)$ where the covariance matrix $\Sigma$ is defined through the squared exponential kernel $K$ by
$\Sigma_{ij}=K(|x_{i}-x_{j}|)=\sigma^{2}\exp\left(-\frac{|x_{i}-x_{j}|^{2}}{2\eta^{2}}\right),$ (32)
where $\sigma^{2}$ is the variance and $\eta$ the length-scale which controls the smoothness. From the quantization rule (31), the likelihood $p(y|f)$ of the observed data is given by
$p(y|f)=\prod_{i=1}^{n}p(y_{i}|f_{i})$
with
$p(y_{i}|f_{i})=\begin{cases}1&\text{if}\quad f_{i}\in[z_{k},z_{k+1}),\\ 0&\text{else.}\end{cases}$
By Bayes’ theorem, the posterior distribution is given by
$\pi(f|y)\propto\exp\left(-\frac{1}{2}f^{\top}\Sigma^{-1}f\right)\bm{1}_{\mathcal{M}}(f),\quad f\in\mathbb{R}^{n}$
where $\mathcal{M}$ is an $n$-dimensional hyper-rectangle determined by the data. Since the covariance matrix $\Sigma$ is highly ill-conditioned due to the smooth kernel $K$, to ensure numerical stability we regularize it as $\Sigma_{\epsilon}=\Sigma+\epsilon I$ for a small $\epsilon>0.$ The barrier function we use here is a linear combination of the corresponding barrier functions in Table 1 for a rectangle when $a_{i}$ and $b_{i}$ are finite and a positive orthant otherwise.
For simulation similarly to [pakman2014exact], we use $n=200,\ \sigma^{2}=0.6,\ \eta^{2}=0.2$ and $\epsilon=10^{-3}$. The locations $(x_{i})_{i=1}^{n}$ are uniformly spaced over the interval $[0,20]$. The quantization thresholds defining the set $\mathcal{M}$ are elements of $\{-\infty,-0.5,0,0.5,+\infty\}$ while the corresponding quantized values $q_{i}\in\{-0.75,-0.25,0.25,0.75\}$. We run WHMC for $10^{6}$ iterations, while MAMLA and MZZS are run for $2\times 10^{6}$ iterations in order to match the same computational time as WHMC. For MZZS, we generate $2\times 10^{6}$ samples from the skeleton and a 10% burn-in is discarded for all methods.
For visualisation, we approximate the true signal using the posterior mean of the samples obtained for each methods and additionally, we quantified the uncertainty by plotting the 25% and 75% posterior quantiles for each of the methods. The results are summarised in Figure 7 below. To further assess the quality of the reconstruction, we show in Figure 8, the evolution of the $L^{2}$-error of the running mean of the posterior samples, using both the WHMC posterior mean ($f_{\text{HMC}}$) and the true signal as references. The results indicate that the reconstruction obtained with MZZS is behaving to a similar fashion to MAMLA.
(a) WHMC
(b) MAMLA
(c) MZZS
(d) Posterior Mean
Figure 7: ayesian reconstruction of a quantized Gaussian process. The bottom-right panel shows the posterior means for each method and the true signal. The remaining panels display the reconstructed signal using WHMC, MAMLA, and MZZS. The shaded region corresponds to the inter-quartile range (25%–75%).
Figure 8: Evolution of the $L^{2}$-error of the running mean of the posterior samples: comparison with the posterior mean calculated using WHMC (left) and with the true function $f$ (right).
5.5 Bayesian logistic regression on constrained domains
In this experiment, we consider a Bayesian logistic regression model with features $x_{j}\in\mathbb{R}^{d}$ and response variable $y_{j}\in\{0,1\}$. We impose a uniform prior on $a,a]^{d}$ which leads to the posterior distribution $\pi$ given by (1) with $\mathcal{M}=[-a,a]^{d}$ and
$U(\theta)=\sum_{j=1}^{n}\log(1+\exp(\theta^{\top}x_{j}))-y_{j}(\theta^{\top}x_{j}).$
Figure 9: A plot of the running relative error of the variance for the two components of the samples generated by MLAm, MZZS, MAMLA, MLAa (left) and SMLAm, SMLAm, MZZ-$S$ (right). Top panel: component $x_{1}$ and bottom panel: component $x_{2}$.
For this experiment we use data given by the German credit dataset [gc] with the features $x_{j}$ being the credit duration and age of the individual and the response $y_{j}$ being the credit outcome of an individual. We have chosen to restrict to $d=2$ as this allows for the calculation of the true posterior expectations by the means of numerical integration. The barrier function for this example is given in Table 1 for the rectangle with $a_{i}=-a,b_{i}=a$ with $a=2$. For this choice of parameter we find that the true posterior variances for each component are $\sigma_{1,\text{true}}^{2}=4.68\times 10^{-3}$ and $\sigma_{2,\text{true}}^{2}=4.31\times 10^{-3}$.
Similarly to Section 5.3 once the samples from each algorithm have been obtained we compute the running relative error of the variance for each component. In addition, we repeat this procedure 50 times and we plot in Figure 9 the averaged relative error for the different algorithms. For sub-sampling algorithms to reduce the noise from using a stochastic gradient, we use a mini-batch of size $31$. We observe that both MZZS and MZZ-SS outperform their SDE-based counterparts, as their relative variance error decreases to zero.
@@ 6 Conclusion and future work
In this paper we proposed a novel class of samplers, mirror PDMPs, which combine PDMPs with the concept of the mirror map. We demonstrated that such samplers benefit from the advantages of PDMPs, namely unbiasedness and the ability to perform exact subsampling while also respecting constraints on the probability distribution. These methods were compared to state of the art SDE-based methods such as PLMC, MYULA, MLAa and MLAm. In our experiments we saw that MZZS and MBPS outperformed the SDE-based algorithms.
@@
In this paper we have concentrated on the most popular PDMPs, ZZS and BPS, however due to the nature of the mirror map it is often the case that the push-forward distribution on the dual space is heavy tailed and therefore it may be beneficial to consider other choices for the deterministic dynamics such as those introduced in [vasdekis2023speed].
Another direction for future work is to include the use of data driven mirror maps such as in [Ye23], this could be combined with an adaptive approach to preconditioning [Bertazzi2020AdaptiveSF]. A final future direction concerns a general issue when working with PDMPs which is to find sharp bounds for the switching rates, this has been addressed in [Corbella22, Bertazzi2022, bertazzi2023splitting, Pagani24] and these techniques can be explored in the context of mirror PDMPs.
@@ Appendix B Calculating Lipschitz constants
In this section, we derive the constants $L_{V},\ L_{V}^{S},\ L_{V}^{L}$ and $M_{\psi}$ which are used to determine the step sizes for the different problems considered in Sections 2.3, 4.1.1 and 5.
@@
Computations of $M_{\psi_{1}}$ and $M_{\psi_{2}}$. Because the functions $x\mapsto-\log(1+x)$ and $x\mapsto-\log(1-x)$ are 1-self concordant, we deduce from [nesterov2018lectures, Theorem 5.1.1] that $M_{\psi_{1}}=\sqrt{2}$ and $M_{\psi_{2}}=2$.
Computation of $L^{L}_{V}$ for $\psi_{1}$ and $\psi_{2}$. For the barrier function $\psi_{1}$, we have for all $x\in[-1,1]^{2}$
$\|\nabla U(x)\|^{2}_{\nabla^{2}\psi_{1}(x)^{-1}}$ $=\frac{(1-x_{1}^{2})^{2}x_{1}^{2}}{1+x_{1}^{2}}+\frac{1}{\epsilon^{2}}\frac{(1-x_{2}^{2})x_{2}^{2}}{1+x_{2}^{2}}$
$\leq\max_{y\in[-1,1]}\frac{(1-y^{2})^{2}y^{2}}{1+y^{2}}\left(1+\frac{1}{\epsilon^{2}}\right)$
$\leq 0.12\left(1+\frac{1}{\epsilon^{2}}\right)=:(L^{L}_{V})^{2}.$
Similarly for the barrier function $\psi_{2}$, we have for all $x\in[-1,1]^{2}$
$\|\nabla U(x)\|^{2}_{\nabla^{2}\psi_{2}(x)^{-1}}$ $=\frac{2(1-x_{1}^{2})^{2}x_{1}^{2}}{1+x_{1}^{2}+(1-x_{1}^{2})^{2}}+\frac{1}{\epsilon^{2}}\frac{2(1-x_{2}^{2})x_{2}^{2}}{1+x_{2}^{2}+\frac{1}{\epsilon}(1-x_{2}^{2})^{2}}$
$\leq\frac{2}{10}\left(1+\frac{1}{\epsilon^{2}}\right)=:(L^{L}_{V})^{2}$
Computation of $L^{S}_{V}$ for $\psi_{1}$ and $\psi_{2}$. We are looking for $L^{S}_{V}$ such that $\nabla^{2}U(x)\prec L^{S}_{V}\nabla^{2}\psi_{i}(x),\ i=1,2.$ For all $v\in\mathbb{R}^{2}$, we have
$\langle\nabla^{2}U(x)v,v\rangle=v_{1}^{2}+\frac{1}{\epsilon}v_{2}^{2}\leq\max\left(1,\frac{1}{\epsilon}\right)\|v\|^{2}_{2}$
$\langle\nabla^{2}\psi_{1}(x)v,v\rangle=\frac{(1+x_{1}^{2})}{(1-x_{1}^{2})^{2}}v_{1}^{2}+\frac{(1+x_{2}^{2})}{(1-x_{2}^{2})^{2}}v_{2}^{2}\geq\|v\|_{2}^{2}.$
Then for the barrier function $\psi_{1}$
$L^{S}_{V}=\max\left(1,\frac{1}{\epsilon}\right).$
Using the fact
$\langle\nabla^{2}\psi_{2}(x)v,v\rangle$ $=\frac{1}{2}\langle\nabla^{2}\psi_{1}(x)v,v\rangle+\frac{1}{2}\langle\Sigma^{-1}v,v\rangle\geq\frac{1}{2}\|v\|_{2}^{2},$
we deduce that for the barrier function $\psi_{2}$, we have
$L^{S}_{V}=\max\left(2,\frac{2}{\epsilon}\right).$
B.2 Calculating $L_{V}$ for MZZS in Section 4.1.1.
Given that
$\nabla\psi_{3}(x)=\left(\frac{x_{1}}{2(1-x_{1}^{2})}+\frac{x_{1}}{2},\frac{x_{2}}{2(1-x_{2}^{2})}+\epsilon^{-1/2}\frac{x_{2}}{2}\right),$
we find $\nabla\psi_{3}^{*}(\zeta)$ by solving the equation $\nabla\psi_{3}(x)=\zeta$ for $x$. This is equivalent to solve the cubic equation
$c_{i}x_{i}^{3}+2\zeta_{i}x_{i}^{2}+(1+c_{i})x_{i}-2\zeta_{i}=0$
for $x_{i}$ as a function of $\zeta_{i}$ where $(c_{1},c_{2})=(1,\epsilon^{-1/2}).$ This will defines $x=\nabla\psi^{*}_{3}(\zeta)$ implicitly. Using the expression of $U$ and the fact that $\nabla\psi_{3}\circ\nabla^{*}_{3}(\zeta)=\zeta,$ we found the expression of $V$ given by
$V(\zeta)=\frac{1}{2}\left(x_{1}^{2}+\frac{x_{2}^{2}}{\epsilon}\right)+\sum_{i=1}^{2}\log\left[\frac{1+x_{i}^{2}}{2(1-x_{i}^{2})^{2}}+\frac{1}{2}c_{i}\right],\quad x=\nabla\psi_{3}^{*}(\zeta).$
The Lipschitz constant $L_{V}$ is therefore given by
$L_{V}=\sup_{\zeta\in\mathbb{R}^{2}}\|\nabla^{2}V(\zeta)\|_{2}=\max_{i=1,2}\sup_{x_{i}\in(-1,1)}\left|\frac{\partial^{2}V}{\partial\zeta_{i}^{2}}\right|$
which we derive to be
$L_{V}=\max\left(1,\frac{4}{(1+\epsilon^{1/2})^{2}}\right).$
B.3 Calculating constants for Section 5.1
Computation of $L_{V}$. For our barrier function, the Fenchel conjugate $\nabla\psi^{*}$ is defined by
@@
Computation of $M_{\psi}$. The barrier $\psi$ is 1-self concordant as the sum of self concordant functions [nesterov2018lectures, Theorem 5.1.1], hence $M_{\psi}=1$.
Computation of $L_{V}^{L}$. For all $x\in(0,+\infty)^{d}$, we have